how do i solve equations with variables on both sides? 7x-11=-19+3x
Answer by nyc_function(2741)

We solve such equations by combining like terms.



7x -11 = -19 + 3x



7x - 3x = -19 + 11


4x = -8


x = -8/4


x = -2